# A Kinetic Model of Solar Wind Acceleration Driven by Ambipolar Electric Potential and Velocity-Space Diffusion

### v1

@@ 6 Conclusions and Future Perspectives
In the exospheric model, the previously introduced Kappa distribution function effectively simulates an increase in the strahl population, thereby accelerating the solar wind to higher velocities. This model does not necessarily require non-Maxwellian distributions at the base of the corona to explain the solar wind acceleration; instead, it relies on a significant strahl population near the Sun. Consequently, the critical question for understanding the solar wind acceleration within the framework of the kinetic exospheric model becomes: How is the strahl formed in the low corona? Further investigations using fully kinetic simulations will be conducted to address this question.

### v2

@@ 6 Conclusions and future perspectives
Furthermore, greater consistency with the current understanding of the diffusion mechanism could be achieved by explicitly distinguishing between the core and escaping populations, as our current approach considers that the same eVDF defines both core and strahl electrons. This refinement could be implemented following the methodology of Zouganelis et al. 2004, in which a Maxwellian distribution and a Kappa distribution were used to define the core and strahl respectively.
In our exospheric model, the previously introduced Kappa distribution function effectively simulates an increase in the strahl population, thereby accelerating the solar wind to higher velocities. However, this model does not necessarily require non-Maxwellian distributions at the base of the corona to explain the solar wind acceleration; instead, it relies on a significant strahl population near the Sun. Consequently, the critical question for understanding the solar wind acceleration within the framework of the kinetic exospheric model becomes: How is the strahl formed in the low corona? Further investigations using fully kinetic simulations will be conducted to address this question.